# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ 3 The Problem of Learning from Imperfect Human Feedback (LIHF)
The user feedback is said to be $\rho$-imperfect for some $\rho\in[0,1]$ if there exists a constant $C_{\kappa}>0$ such that the corruption $c_{t}$ satisfies $|c_{t}|\leq C_{\kappa}t^{\rho-1},\forall t\in[T]$. [^1] All our results naturally applies to $\rho<0$, which is a significantly easier situation since accumulated corruption is a constant in that case. We hence will not explicitly consider it in this paper.
For a $\rho$-imperfect user, the accumulated corruption $\sum_{\tau=1}^{t-1}|c_{\tau}|\leq C_{\kappa}t^{\rho}$ for every $t$, which is why this is coined $\rho$-imperfect. The diminishing inaccuracy of user estimation in Definition 1 is a natural model for user’s learning behavior and has already been used in various previous works to model user’s imperfect feedback (Wang et al., 2023; Yao et al., 2022a; Immorlica et al., 2020) (Details are presented at Appendix A). Note that agnostic corruption in this situation means the learner know $|c_{t}|\leq C_{\kappa}t^{\rho-1}$ for some $\rho$, but she does not know the value of $\rho$.
On the other hand, the corruption are said to be arbitrary up to certain budget $C$, if $\sum_{t=1}^{T}|c_{t}|\leq C$ for a given (hence fixed) time horizon $T$. Our designed (positive) learning algorithmic will have guaranteed regret upper bound under arbitrary corruption, though our lower bound (i.e., negative result) will hold even in the weaker $\rho$-imperfect user feedback.
We remark that another line of recent research studies a different type of corruption that directly flip the outcomes of the duels; i.e., flipping the realized event $a\succ a^{\prime}$ to its opposite $a^{\prime}\succ a$ (Di et al., 2024; Saha and Gaillard, 2022; Komiyama et al., 2015). These corruption models are motivated by the existence of true adversaries, e.g., malicious users in recommendation systems (Lykouris et al., 2018) or fraudulent clicks (Deshpande and Montanari, 2013), who have the capability of altering user choices. In contrast, only utilities are corrupted in our model. These two models are not comparable. For instance, the adversary can induce a deterministic event $a\succ a^{\prime}$ at any given round with one corruption, whereas the outcomes in our model is almost always probabilistic (unless the adversary uses an infinite budget $c_{t}$). Such modeling difference inherently comes from the different motivations of our LIHF problem. Specifically, our model is to capture imperfect human feedback due to inaccurate perception of utilities or ignorance of true preferences. For example, when recommender systems learn to find the optimal recommendation for a user, the user may not know her true preference at the beginning and often gradually refines her preferences through experiencing different items (Yao et al., 2022b). Similarly, when generative models like ChatGPT learn to generate personalized content for a user, the user may go through a similar procedure of preference refinement during interactions with the model. These modelling differences also necessitate our fundamentally different techniques. Our approaches are based on gradient-based methods whereas Di et al. (2024) and Saha (2021) employ the principle of optimistic in face of uncertainty and construct the confidence set of preferences through querying the maximum information pair. Their approaches inherently require the knowledge of total corruption level $C$. In contrast, the analysis of our gradient-based methods controls the impact of gradient bias (due to corruption) on regret. By adjusting its learning rate, our algorithm will gracefully “trade” its learning efficiency for more robustness against agnostic corruption. Moreover, our algorithm is suitable for infinite action space, whereas previous works of Di et al. (2024); Saha (2021) are for finite action spaces.
@@
Learning goal: agnostic corruption-robust regret minimization. The goal of the learner is to optimize her sequential actions to minimize the following dueling regret against an arbitrary sequence of corruption $\{c_{t}\}_{t=1}^{T}$:
$\text{Reg}_{T}=\mathbb{E}\left(\sum^{T}_{t=1}\sigma(\mu(a^{*})-\mu(a_{t}))+\sigma(\mu(a^{*})-\mu(a^{\prime}_{t}))-1\right).$ (1)
This regret measure is also studied in many prior literature (Saha and Gaillard, 2022; Kumagai, 2017; Komiyama et al., 2015; Yue and Joachims, 2009). Importantly, all our algorithms are agnostic to the corruption in the sense that the algorithm does not depend on any knowledge of $\{c_{t}\}_{t=1}^{T}$.
4 The Intrinsic Limit of Learning from Imperfect Human Feedback
To understand the inherent difficulty of LIHF, in this section we develop regret lower bounds for a setting that is even easier to learn – that is, learning from a $\rho$-imperfect user (Definition 1) with a known $\rho$. While this non-agnostic may appear simpler, our main result of this section exhibits a tight lower bound of $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ for strongly concave user utilities. We also show a similar lower bound hold for linear user utility as well. Our lower bound result differs in two key aspects from various previous lower bound results for online learning under corruptions (Agarwal et al., 2021; Bogunovic et al., 2020). Firstly, previous lower bound proofs require the adversaries’ capability to corrupt the reward feedback in an arbitrary way, whereas the corruption in our proof will follow Definition 1. It is unclear whether such strictly more restrictive corruption will be more tractable since it does offer the learner much more accurate estimation about the amount of corruption at each round. Secondly, previous construction of hard instances that attains regret lower bound relies on linear reward, whereas our lower bound holds for strongly concave utility functions, since they more realistically model human’s risk attitude and consumption utilities (Chiappori and Rochet, 1987). Strongly concave utility functions are generally perceived as easier to optimize and to learn, hence it is unclear whether it is also the case under imperfect user feedback. While both differences above seem to hint that our problem might be easier than previous online learning problem under corruption, we show that the answer is actually “No”.
Theorem 1 (A Tight Lower Bound).
There exists an LIHF instance with a $\rho$-imperfect user (Def. 1) and strongly concave user utility such that any learner has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$ even with the knowledge of $\rho$. Moreover, this regret lower bound is tight, up to logarithmic terms. That is, for the same setting, there exists an online learning algorithm with $\text{Reg}_{T}\leq\tilde{O}(d\max\{\sqrt{T},T^{\rho}\})$.
This theorem reveals the fundamental limits of learning from $\rho$-imperfect users, even when the learner knows $\rho$. The core new lemma needed for proving the regret lower bound in Theorem 1 is to show that even in standard online learning with noisy bandit feedback about any pulled arm’s reward (as opposed to comparative feedback in dueling setups), the same regret lower bound holds. This is summarized in the following Lemma 1. We then convert this lower bound for standard bandit reward feedback to our setting of dueling/comparative feedback. Notably, Ye et al. (2024b) recently showed a matching upper bound for the standard bandit reward feedback setting assuming arbitrary corruption with total amount $C$. Hence Lemma 1 and their results together echoes our key insight in this section — that is, learning from $\rho$-imperfect user is as hard as learning from arbitrarily corrupted rewards in classic bandit reward feedback setting as well.
For linear user utility, perhaps unsurprisingly, we are able to show the same lower bound as follows (see Prop. 1). En route, we similarly show that the problem of learning directly from bandit linear reward feedback also suffers the same regret lower bound $\Omega(d\max\{\sqrt{T},T^{\rho}\})$. This happens to resolve an open problem proposed by Yao et al. (2022a) and implies that the regret upper bound shown in (Yao et al., 2022a) for the “learning from learning user” problem is tight in $T$.
@@ 4 The Intrinsic Limit of Learning from Imperfect Human Feedback
There exists an LIHF instance with a $\rho$-imperfect user and linear user utility such that any learner knowing $\rho$ has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$.
We defer the full proof of Theorem 1 and Proposition 1 to Appendix B.1 and B.2. Below we only state a representative core lemma for the strongly concave utility function case and its proof sketch, which is the crux of our lower bound proof. Notably, the proof for the matching upper bound in Theorem 1 is also highly non-trivial. Our starting point is the recent NC-SMD algorithm. However, to achieve the tight regret upper bound, we have to employ a novel and highly involved induction argument to identify the optimal learning rate. Through this process, we established a new connection between the convergence of the $L_{2}$-norm of the action $\|a_{t}-a^{*}\|_{2}$ and the regret itself (It has been experimentally verified at Appendix D.1.2).

### v2

@@ 2 The Problem of Learning from Imperfect Human Feedback
The human feedback is said to be $\rho$-imperfect for some $\rho\in[0,1]$ if there exists a constant $C_{\kappa}$ such that corruption $c_{t}(a_{t},a^{\prime}_{t})$ satisfies $|c_{t}(a_{t},a^{\prime}_{t})|\leq C_{\kappa}t^{\rho-1},\forall t\in[T]$. [^2] All our results naturally applies to $\rho<0$, which is a significantly easier situation since accumulated corruption is a constant in that case. Therefore, we will not explicitly consider it in this paper.
The definition shows that $\rho$-imperfect human feedback is mathematically equivalent to a restricted form of strong adversarial corruption to user utility functions. The total corruption is bounded by $\sum_{t=1}^{T}|c_{t}(a_{t},a^{\prime}_{t})|\leq C_{\kappa}T^{\rho}$, which is a key parameter influencing the intrinsic difficulty of the learning problem. We thus cast the Learning from Imperfect Human Feedback (LIHF) problem as dueling bandits under strong corruption but with decaying scales. For clarity, we use “arbitrary adversarial corruption” as corruption without decaying constraints and “$\rho$-Imperfect Human Feedback” as corruption with decaying constraints, as described in Definition 1. Through this modeling, we aim to: first create a more optimistic framework for LIHF compared to arbitrary adversarial corruption, and second, explore LIHF’s intrinsic difficulty to achieve more efficient learning guarantees than those possible under arbitrary adversarial corruption.
Several points are worth noting about Definition 1. While we are not the first to model imperfect human feedback, we are the first to apply it to dueling bandits with strictly concave utilities. Recent works, motivated by recommendation systems, have explored learning user preferences from imperfect feedback (Immorlica et al., 2020; Yao et al., 2022; Wang et al., 2023). These studies assume that users do not know their true expected reward $\theta_{i}$ for each choice $i$ (i.e., arm) and instead behave based on an estimated reward $\hat{\theta}_{t,i}$ at time $t$. The utility difference $|\theta_{i}-\hat{\theta}_{t,i}|=c_{t}$ is modeled as a decreasing function of time, $t^{\rho-1}$, indicating that users improve their preference estimates over time. By viewing the human utility $\mu(a)$ as the average reward, our modeling of $\rho$-Imperfect Human Feedback shares the same spirit; it captures imperfect yet gradually improving human feedback (e.g., due to inaccuracies in utility perception or initial ignorance of true preferences). As one of our motivations, when generative models like ChatGPT learn to create personalized content, the user could undergo a process of preference refinement during interactions with the model. Additionally, our model differs from recent corruption-robust dueling bandit studies (Komiyama et al., 2015; Saha and Gaillard, 2022; Di et al., 2024), which corrupt realized outcomes (e.g., flipping $a^{\prime}\succ a$ to $a\succ a^{\prime}$) and are motivated by adversaries like malicious users or fraudulent clicks (Deshpande and Montanari, 2013; Lykouris et al., 2018). In contrast, our model focuses on utility corruption caused by imperfect perception of true utility, not outcome manipulation.
@@
Learning goal: regret minimization by Learning from $\rho$-Imperfect Human Feedback ($\rho$-LIHF). The goal of the learner is to optimize her sequential actions to minimize the following dueling regret in the $\rho$-LIHF problem:
$\text{Reg}_{T}:=\mathbb{E}\left(\sum^{T}_{t=1}\sigma(\mu(a^{*})-\mu(a_{t}))+\sigma(\mu(a^{*})-\mu(a^{\prime}_{t}))-1\right).$ (1)
This regret measure has been extensively studied in prior literature (Yue and Joachims, 2009; Komiyama et al., 2015; Kumagai, 2017; Saha and Gaillard, 2022). Other regret measures, such as the cumulative difference between optimal and obtained utility, are also considered. We discuss the equivalence of various regret measures in Lemma 15.
3 The Intrinsic Limit of LIHF
In this section, we study the intrinsic limit of learning from $\rho$-Imperfect Human Feedback. We are particularly interested in understanding whether learning from this restricted version of corrupted feedback is fundamentally “easier” than learning from arbitrarily adversarial corruption. Somewhat surprisingly, the answer seems to be no, as illustrated by the following lower bound result.
Theorem 1.
There exists a $\rho$-Imperfect Human Feedback (see Definition 1), strongly concave utility function $\mu$, and link function $\sigma$ such that any learner has to suffer $\text{Reg}_{T}\geq\Omega\left(d\max\{\sqrt{T},T^{\rho}\}\right)$, even with the knowledge of $\rho$.
Notably, similar lower bound results have appeared in recent studies on adversarial corruption in bandits (Bogunovic et al., 2022), including dueling bandits (Agarwal et al., 2021; Di et al., 2024), showing the same lower bound $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ (though they often use $C:=\Theta(T^{\rho})$ to denote total corruption budget). However, a few key distinctions are worth noting. First, Theorem 1 presents a stronger result, showing that even when corruption decays over time, this structure (which defines our LIHF problem) does not make learning easier. Second, our corruption targets utility values, which is quite different from the corruption of comparison outcomes studied in (Agarwal et al., 2021; Di et al., 2024). In their case, corruption can fully flip comparison outcomes, which is critical for promoting sub-optimal arms in their lower bound proofs. However, such full control of the outcome is infeasible in our corruption of only the utility, which only slightly shifts the comparison probabilities since the scale of the corruption is also upper bounded by $\mathcal{O}(t^{\rho-1})$.
The aforementioned difference from previous works (Agarwal et al., 2021; Di et al., 2024) also render our proof of Theorem 1 fundamentally different from (and significantly more involved than) their proofs. Specifically, due to limited and diminishing corruption power, the corruption in our LIHF problem are insufficient to completely “mask” the instance as one with a different sub-optimal arm, which is the key strategy in previous lower bound proofs. Thus, our proof has to leverage information-theoretic lower bounds of statistical distributions to understand how small-scale corruption at each round could influence the overall function estimation errors. At the core of our proof is the following lower bound result for a different, and intuitively easier, problem: the standard bandit setup with direct reward feedback, as shown in Lemma 1 below. We then convert this lower bound for direct reward feedback to dueling feedback through a linear link function, i.e. $\sigma(x)=\frac{1+x}{2}$.
Lemma 1 (Lower Bound under Direct Reward Feedback).
Consider bandit learning with direct reward feedback, where reward $r(a):=\mu(a)+\epsilon$, $\epsilon$ follows standard normal distribution. The action space $\mathcal{A}$ is contained in a $d$-dimensional unit ball, and $\mu$ is $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, with $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$ Under $\rho$-Imperfect Human Feedback (Definition 1), for any $T$ and $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, there exists a $\theta$ such that for any learner under direct reward feedback, even with the knowledge of $\rho$ [^3] We only consider $\rho\geq\frac{1}{2}$. If $\rho<\frac{1}{2}$, the lower bound degenerates to $\Omega(d\sqrt{T})$ (Kumagai, 2017)., has to suffer regret $\text{Reg}_{T}:=\mathbb{E}\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\}\geq\frac{d}{4}C_{\kappa}T^{\rho}.$
Proof [Proof Sketch of Lemma 1] We aim to show that there exists a corruption strategy such that the regret incurred by any learner with direct reward feedback is no less than $\Omega(dT^{\rho})$. The formulation of such a corruption strategy hinges on the observation that the regret incurred by the learner is bounded below by the sum of the Kullback-Leibler (KL) divergence between the distribution of the corrupted reward feedback $\hat{v}_{t}$ conditioned on different values of $\theta$. Specifically, let $\theta$ be uniformly drawn from $\{-\beta,\beta\}^{d}$, for some constant $\beta>0$. Given an action $a_{t}$ and index $i$, conditioned on the value of the $i$-th coordinate of $\theta$, $\theta_{i}$, if $\theta_{i}>0$, then $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}>0)+\xi_{a_{t}}=\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)+\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}>0)+\xi.$
Conditioned on $\theta_{i}<0$, the corrupted reward feedback $\hat{v}^{\prime}_{t}$ is
$\hat{v}^{\prime}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}<0)+\xi_{a_{t}}=\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)-\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}<0)+\xi.$
The noise $\xi$ follows standard normal distribution. Consider the following corruption strategy which sets $c_{t}(a_{t}|\theta_{i}>0):=-\beta a_{t,i}$ and $c_{t}(a_{t}|\theta_{i}<0):=\beta a_{t,i}$ to minimizes the KL divergence between $\hat{v}_{t}$ and $\hat{v}^{\prime}_{t}$. Together with $1$-strongly concavity of $\mu_{\theta}$, we can establish $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq\frac{1}{2}\max\left\{\sum^{T}_{t=1}(\frac{C_{\kappa}T^{\rho-1}}{\beta}-\beta)^{2},\frac{dT\beta^{2}}{2}\right\}.$ Since $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, we can set $\beta:=\sqrt{C_{\kappa}}T^{\frac{\rho-1}{2}}$ to satisfy $\|\theta\|_{2}\leq 1$ and to optimize the aforementioned lower bound. Then we obtain $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq dC_{\kappa}T^{\rho}/4$. Notice that the corruption level each round is bounded by $\beta\|a_{t}\|_{\infty}$. In the scenario when the radius of $\mathcal{A}$ is upper bounded by $\beta$, the corruption budget each round is bounded by $C_{\kappa}t^{\rho-1}$. It implies that the corruption is $\rho$-Imperfect Human Feedback, completing the proof.
Theorem 1 can be proved by converting the lower bound in Lemma 1 for direct reward feedback to dueling feedback (see Lemma 7). Before concluding this section, we note that the strongly concave utility function in the lower bound of Theorem 1 is not necessary. In particular, Proposition 1 below shows that similar lower bound applies to linear utilities as well. This result happens to resolve an open problem posed by Yao et al. (2022) which studies a similar learning from imperfect user problem with linear utility and dueling feedback (they termed it the “learning from a learning user” problem), develops a no-regret learning algorithm for that setting, but left the lower bound as an open problem. Proposition 1 implies that their upper bound’s dependence on $T$ is tight. We direct readers to Appendix A.1 and A.2 for proof details of Lemma 1 and Proposition 1.
@@ 3 The Intrinsic Limit of LIHF
There exists an LIHF instance with $\rho$-Imperfect Human Feedback (Definition 1), linear user utility $\mu$, and link function $\sigma$, such that any learner has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, even with the knowledge of $\rho$.
4 An Efficient and Tight $\rho$-LIHF Algorithm
In this section, we present a learning algorithm with a regret upper bound that matches the $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ lower bound, up to a logarithmic term, for the $\rho$-LIHF setting discussed in Section 3. Our algorithm, Robustified Stochastic Mirror Descent for Imperfect Dueling (RoSMID), outlined in Algorithm 1, generalizes the Noisy Comparison-Based Stochastic Mirror Descent (NC-SMD) from Kumagai (2017). RoSMID employs a corruption-aware learning rate, $\eta_{\rho}:=\sqrt{\log T}/(dT^{\max\{0.5,\rho\}})$, to slow down the learning when the imperfection level $\rho$ is large, with NC-SMD being the special case with $\rho=0$. Although RoSMID may initially seem like a pretty natural extension of NC-SMD, the choice of optimal $\eta_{\rho}$ is in fact through careful derivation from a novel framework for analyzing continuous bandits under utility corruption. We conclude by demonstrating the applicability of this new framework to analyzing other gradient-based algorithms.
Theorem 2.
RoSMID satisfies $\text{Reg}_{T}\leq\tilde{\mathcal{O}}(d\max\{\sqrt{T},T^{\rho}\})$ for any $\rho$-LIHF problem.
The main technical contribution underlying this theorem is a novel framework for analyzing continuous dueling bandits with arbitrary strongly concave, yet possibly imperfect or corrupted utilities, which we believe has independent value. This framework not only enables a tight regret upper bound analysis, as shown in Theorem 2, but also allows us to easily derive regret guarantees for other dueling bandit algorithms under adversarial corruption, as demonstrated in Subsection 4.1. While recent works have explored corruption-robust regret analysis for dueling bandits with $K$ arms (Saha and Gaillard, 2022) or linear utilities (Di et al., 2024), neither their algorithmic approaches nor their analysis techniques can be applied to our setting with arbitrary concave utilities in continuous action space.
Input: Learning rate $\eta_{\rho}:=\sqrt{\log T}/\left(dT^{\max\{0.5,\rho\}}\right)$, $\nu$-self-concordant function $\mathcal{R}$  [^4] We direct reader to Definition 2 for formal definition of self-concordancy., time horizon $T$, tuning parameters $\lambda\leq\frac{l^{\sigma}_{2}\alpha}{2}$, $\phi\geq((L^{\sigma}_{1})^{3}L^{\sigma}_{2}/\lambda)^{2}$
1 Initialize $a_{1}=\arg\min_{a\in A}\mathcal{R}(a)$
2 for $t\in[T]$ do
3       Update concordant function $\mathcal{R}_{t}(a)=\mathcal{R}(a)+\frac{\lambda\eta_{\rho}}{2}\sum^{t}_{i=1}\|a-a_{i}\|^{2}+\phi\|a\|^{2}$
4       Sample direction $u_{t}$ uniformly at random from a unit sphere $\mathbb{S}^{d}$
5       Obtain corrupted feedback ^$\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})$$\in\{0,1\}$, for $a_{t}$ and $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-1/2}u_{t}$
6       Compute the corrupted gradient: $\hat{g}_{t}=\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})d\cdot\nabla^{2}\mathcal{R}_{t}(a_{t})^{1/2}\cdot u_{t}$
Output : $a_{T+1}$
Algorithm 1 Robustified Stochastic Mirror Descent for Imperfect Dueling (RoSMID)
The full proof of Theorem 2 is quite involved and is deferred to Appendix B. Here, we outline the main ideas, divided into four key steps. Step 1-3 are applicable for analyzing continuous dueling bandits under any form of corruption, whereas last Step 4 is the only step that hinges on the decaying corruption level assumptions of $\rho$-LIHF and is also the most involved and novel step.
Proof [Proof Sketch] The proof of Theorem 2 is divided into four major steps.
Step 1: quantifying bias of gradient estimation in $\rho$-LIHF
Our proof starts from quantifying the bias of the gradient $\hat{g}_{t}$ caused by $\rho$-Imperfect Human Feedback, as shown in the following Lemma 2.
Lemma 2 (Corrupted Gradients Estimation).
The gradient $\hat{g}_{t}$ in Line 6 of Algorithm 1 satisfies
$\mathbb{E}\left(\hat{g}_{t}|a_{t}\right)=\nabla|_{a=a_{t}}\bar{P}_{t}(a)-\mathbb{E}\left(b_{t}|a_{t}\right).$
where $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[\mathbb{P}(a_{t}\succ a+\nabla^{2}R_{t}(a_{t})^{-\frac{1}{2}}x)\right]$ (with $\mathbb{B}$ as the unit ball) is the smoothed probability that $a_{t}$ is preferred over any action $a$ , and $b_{t}=d\left(\mathbb{P}(a_{t}\succ a^{\prime}_{t})-\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})\right)\nabla^{2}\mathcal{R}_{t}(a_{t})^{\frac{1}{2}}u_{t}$.
That is, the scale of bias at round $t$, term $b_{t}$, defined in Lemma 2, depends linearly on $a$’s dimension $d$ and the difference between the true and corrupted probability that $a_{t}$ is preferred over $a^{\prime}_{t}$. The direction of $b_{t}$ is determined by the sampled direction $u_{t}$ projected by the square root of Hessian of the self concordant function $\mathcal{R}_{t}$, evaluated at $a_{t}$. This characterization is essential for our following regret decomposition.
Step 2: decomposing regret for dueling bandits into regret of decision and feedback error
Core to our proof is a “regret decomposition” lemma below that upper bounds the regret by two parts: regret of sub-optimal decision making under uncertainty, referred as regret of decision (first term in (2)), and regret due to corruption, referred as feedback error (second term in (2)).
Lemma 3 (Regret Decomposition for Dueling Bandits).
The $\text{Reg}_{T}$ of Algorithm 1 satisfies:
$\text{Reg}_{T}\leq\underbrace{\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Regret\hskip 2.84526ptof\hskip 2.84526ptDecision}}+\underbrace{2\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Feedback\hskip 2.84526ptError}},$ (2)
where $C_{0}:=(2\nu+4L^{\sigma}_{1}\kappa+4R^{\sigma}_{2}(L^{\mu})^{2}+L^{\mu}L^{\sigma}_{1}\kappa)/\lambda$, and $a^{*}_{T}:=\frac{1}{T}a_{1}+(1-\frac{1}{T})a^{*}$.
We remark that similar regret decomposition has been shown to be useful for reinforcement learning which has direct reward feedback (Foster et al., 2023). However, to the best of our knowledge, Lemma 3 is the first to exhibit such a decomposition for dueling feedback, which contains much sparser information than direct reward feedback as in classic RL or online learning. Hence the lemma may be of independent interest for future research on dueling bandits under corruption or erroneous feedback.
Step 3: upper bounding the regret from feedback error
Bounding the regret of decision term in (2) is standard. Thus, this step develops a novel upper bound for bounding the regret from feedback error, as shown below.
Lemma 4 (Feedback Error).
The feedback error term in (2) can be bounded as follows:
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq C_{1}d\sqrt{\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|\sqrt{\mu(a^{*})-\mu(a_{t})}\right\}+C_{2}dT^{\rho}+2d\sqrt{\lambda}\sqrt{\eta_{\rho}T},$
where $C_{1}:=\sqrt{\frac{2\lambda(L^{\sigma})^{2}}{\alpha}}$, and $C_{2}:=2L^{\sigma}R\sqrt{\sup_{a\in\mathcal{A}}\lambda_{\max}\left(\nabla^{2}\mathcal{R}(a)\right)+2\phi}$.
To our knowledge, such a bound is new for dueling bandits. The less common term in the bound is $\sqrt{\mu(a^{*})-\mu(a_{t})}$. It comes from the strong concavity of the utility function $\mu$, which implies $\langle b_{t},a_{t}-a^{*}\rangle\leq\|b_{t}\|_{2}\|a_{t}-a^{*}\|_{2}\leq\|b_{t}\|_{2}\sqrt{\frac{2}{\alpha}(\mu(a^{*})-\mu(a_{t}))}$.
Step 4: The iterative regret refinement analysis and resultant optimal learning rate
Our last but the most intriguing step is to pin down the regret upper bound. Given the bounds from previous steps, the only remaining term to upper bound is
$\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|\sqrt{\mu(a^{*})-\mu(a_{t})}\right\}$ in Lemma 4. We can upper bound $\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|$ by $C_{\kappa}t^{\rho-\frac{1}{2}}$ under $\rho$-Imperfect Human Feedback assumption (this is the point where we start to need the decaying structure in Definition 1). This is a tricky question, since this term involving optimal action $a^{*}$ closely connects to the regret $\text{Reg}_{T}$ itself (see (1)). This is precisely the motivation for our analysis approach of “iterative regret refinement”. Concretely, it turns out that a good upper bound for the term $A:=\mathbb{E}\left(\sum^{T}_{t=1}\mu(a^{*})-\mu(a_{t})\right)$ can be leveraged to derive a good upper bound for term $B:=\mathbb{E}\left\{\sum^{T}_{t=1}t^{\rho-\frac{1}{2}}\sqrt{\mu(a^{*})-\mu(a_{t})}\right\}$, which can then be leveraged to derive a good upper bound for the regret $\text{Reg}_{T}$, which can be used to refine the original upper bound of term $A$, at which point we can re-apply the above refinement process.
Two key factors are essential for the iterative regret refinement mentioned above to result in a tight bound. The first step is to derive a good upper bound for term $B$ using term $A$, which is shown in Lemma 5. The second step is to prove that this iterative refinement consistently improves the regret bound, eventually reaching a limit. This is demonstrated by Lemma 6. The choice of learning rate $\eta_{\rho}=\frac{\sqrt{\log T}}{dT^{\max\{1/2,\rho\}}}$ is crucial in the second step; in a sense, it is the only choice that converges to the tight regret upper bound $\tilde{\mathcal{O}}\left(d\max\{\sqrt{T},T^{\rho}\}\right)$ in the limit, rather than exploding the bound.